Solve the following system of linear equations using the substitution method:
3x-2y=18
5x+10y=-10
Express answer as an ordered pair

Answer by Alan3354(69443)

Solve the following system of linear equations using the substitution method:
3x-2y=18
5x+10y=-10--> x + 2y = -2 --> x = -2y - 2
---
Sub for x in the 1st eqn
3(-2y-2) - 2y = 18
-6y - 6 - 2y = 18
-8y = 24
y = -3
===========
etc